OpenStudy (anonymous):

f(x)=2x-2/x^2+x-2 What values of x is f(x) discontinuous?

OpenStudy (anonymous):

\[f(x)=\frac{ 2x-2 }{ x^2+x-2 }\]

OpenStudy (anonymous):

I've done the work and ended up with \[\frac{ 2 }{ x+2 }\] but I do not know about the discontinuities.

OpenStudy (anonymous):

a=f(x) =(2x-2)/(x^2+x-2) =(2x-2)/(x-1)(x+2) =2/(x+2) f(x) is undefined at x= -2 & x=1 b=lim 2/(x+2) = -infinity x-->-2 c)x=-2 lim 2/(x+2) = 0 x->infinity d)2/(x+2) = a/(b + x)

OpenStudy (anonymous):

....what..

OpenStudy (anonymous):

f(x) is undefined at x= -2 & x=1

OpenStudy (anonymous):

That's not what the problem is asking for. It's asking me to define the discontinuities. I'm aware that x=-2 and x=1 are non-removable discontinuities but I don't know if I'm missing anything else.

OpenStudy (anonymous):

. . . Seems like you got it.

OpenStudy (anonymous):

Thanks love. :)

OpenStudy (anonymous):

Mhm. ^^